OpenStudy (anonymous):

The registrar of a certain university estimates that the total student enrollment in the Continuing Education division will be given by N(t)=-(20,000/square root(1+0.01t)) + 21000 where N(t) denotes the number of students enrolled in the division t years from now. Find an expression for N'(t). I got 100(1+0.01t)^-3/2 and it said that was wrong so i tried 100/(1+0.01t)^3/2 and it said was wrong too

OpenStudy (anonymous):

wish i could read this

OpenStudy (anonymous):

well its only the 1+0.01t thats squared i don't have a fancy keyboard to make that anymore readable then i have it kinda stinks

OpenStudy (anonymous):

(20000/squared(1+0.01t)) +21000

OpenStudy (dumbcow):

is this it \[N(t) = 21000 - \frac{20000}{(1+.01t)^2}\]

OpenStudy (anonymous):

i mean the square root of 1 + 0.01t im sorry let me change it

OpenStudy (dumbcow):

\[N(t) = 21000 - \frac{20000}{\sqrt{1+.01t}}\]

OpenStudy (dumbcow):

\[N'(t) = 21,000 - (-\frac{1}{2})(.01) \frac{20,000}{(1+.01t)^{3/2}}\]

OpenStudy (anonymous):

yes

OpenStudy (dumbcow):

sorry the 21,000 goes away you are left with \[\frac{100}{(1+.01t)^{3/2}}\]

OpenStudy (anonymous):

yeah thats what i got and it said it was wrong

OpenStudy (anonymous):

i actually have that as one of my answers i got on top i tried formatting it a different way and all

OpenStudy (dumbcow):

ahh i see, well lets check... http://www.wolframalpha.com/input/?i=differentiate+21%2C000+-+20%2C000%2Fsqrt%281%2B.01t%29 its correct, maybe check you didn't write the original function incorrectly

OpenStudy (anonymous):

nope wrote it all right and everything it's just web assign i have to put in in a simplified or less simplified way actually how it wants it, but it just doesn't ever tell you how it wants it so your left guessing theres been a lot of complaints about it though

OpenStudy (dumbcow):

oh ok, try writing it in radical form \[\frac{100}{\sqrt{(1+.01t)^3}}\]

OpenStudy (anonymous):

it still won't take it and i don't think i can submit anymore its ok thank you though

OpenStudy (dumbcow):

yw , well at least you know you did the math correctly :)